# Few-shot Adaption to Distribution Shifts By Mixing Source and Target Embeddings

### v2

@@ 2 Related Work
Few-shot adaption and sample efficiency. In scenarios where labeled target data are available but scarce, adapting the last layer to the target distribution in a data-efficient manner presents a significant challenge. The study most relevant to our paper is (Chen et al., 2023), which considers the same setting as ours, where the number of available target data for linear probing is very small (e.g., 2 to 32 per class). They propose Pro2 that first learns a projection that maps the source data embeddings to orthogonal informative directions. Then, it passes the target data embeddings through the learned projection and perform linear probing using these projected embeddings. (Teney et al., 2022)’s method shares a similar intuition but does so with an additional loss term instead of explicitly enforcing orthogonality. However, we find that in both methods, the projection layer learned on the source data may miss directions in the embeddings that are important for the target distribution, as we will theoretically demonstrate. In contrast, our method does not suffer from this issue. We include these two methods as baselines in our experiments.
@@ 5.1 A case study of domain generalization: the advantage of MixPro over prior work
[…]
@@ 5.2 A case study of subpopulation shift: trade-off between learning target and overfitting noise
[…]
@@ 6 Experiments
Datasets. We consider 8 datasets with different types of distribution shifts. These include 3 subpopulation shift benchmarks: WaterBirds (Sagawa et al., 2019), bFFHQ (Kim et al., 2021), and UrbanCars (Li et al., 2023); along with 5 domain generalization benchmarks: PACS (Li et al., 2017), VLCS (Fang et al., 2013), Office-Home (Venkateswara et al., 2017), Terra Incognita (Beery et al., 2018), and Camelyon17 (Koh et al., 2021). Due to space limitations, we defer the details to Appendices B.1 and B.2. For the target dataset size, we consider the few-shot scenario where a range of {2, 4, 8, 16} data points per class are sampled from the target distribution. The test set is constructed using the remaining data in the target distribution.
@@ A.2.1 Analysis of MixPro
By applying the classical result for the inverse of perturbation (Demmel, 1992) and combining it with 18, we obtain
@@
$\bm{u}_{\text{mixed}}\coloneqq\mathbbm{E}\bm{u}_{\text{mixed}}=\begin{bmatrix}\frac{1}{2}(1-s)\\ \frac{1}{2}\\ \frac{1}{2}s\\ \end{bmatrix}.$ (23)
@@
$\bm{\Sigma_{t}}=\begin{bmatrix}\sigma^{2}&0&0\\ 0&\frac{1}{2}&0\\ 0&0&\frac{1}{2}\end{bmatrix}.$ (32)
@@ A.2.2 Analysis of Pro2
$=$ $\bm{\Pi}_{i}^{\top}\begin{bmatrix}1/2&0&0\\ 0&1/2&0\\ 0&0&\sigma^{2}\end{bmatrix}\bm{\Pi}_{i}-[1,~{}~{}1,~{}~{}0]~{}~{}\bm{\Pi}_{i}+1.$ (36)
@@
$\bm{\Pi}_{i}^{*}=\arg\min_{\bm{\Pi}_{i}}\bigg{(}\bm{\Pi}_{i}^{\top}\begin{bmatrix}1/2&0&0\\ 0&1/2&0\\ 0&0&\sigma^{2}\end{bmatrix}\bm{\Pi}_{i}+1\bigg{)}.$ (40)
@@
$\hat{\mathbbm{E}}_{(\bm{z}^{t},y)\in\mathcal{E}^{t}}(\bm{\Pi}^{*\top}\bm{z}^{t}\bm{z}^{t\top}\bm{\Pi}^{*})=\hat{\mathbbm{E}}_{(\bm{z}^{t},y)\in\mathcal{E}^{t}}\begin{bmatrix}(\bm{z}^{t\top}\bm{\Pi}_{1}^{*})^{2}&0&0\\ 0&0&0\\ 0&0&0\end{bmatrix}$ (45)
@@ A.3 Effect of $s$ when $\sigma=\omega(1)$
$\mathbbm{E}_{(\bm{z},y)\in\mathcal{E}^{t}}(\bm{w}_{MixPro}^{*\top}\bm{z}-y)^{2}=$ $\bigg{(}\frac{q}{\psi_{1}\psi_{3}-\psi_{2}^{2}}(\psi_{3}-\psi_{2}(1-s)(2p-1))-1\bigg{)}^{2}+\bigg{(}\frac{q}{\psi_{1}\psi_{3}-\psi_{2}^{2}}(-\psi_{2}+\psi_{1}(1-s)(2p-1))\bigg{)}^{2}$
@@
$\frac{1}{n}\bm{Z}\bm{Z}^{\top}=\frac{1}{n}\sum_{i=1}^{n}\bm{z}\bm{z}^{\top}=\begin{bmatrix}\bm{G}_{spu}&\bm{0}\\ \bm{0}&\frac{\sigma_{\xi}^{2}}{d-2}\bm{I}_{d-2}\end{bmatrix},\text{ where }~{}~{}\bm{G}_{spu}\coloneqq\begin{bmatrix}1+\sigma_{1}^{2}&2p-1\\ 2p-1&1\end{bmatrix}$ (55)
@@
The following derivations are performed based on the equations in Proposition A.1. We index the examples such that the first $m/2$ examples in $\mathcal{E}^{t}$ have label $1$ and the rest have label $-1$. Since equations 49 and 50 are true, without loss of generality, we assume $\bm{\zeta}^{\prime}_{i}=\sigma_{\xi}\bm{e}_{i}$ where $\bm{e}_{i}$ is the $i$-th standard basis in $\mathbbm{R}^{d-2}$.
@@
$\frac{1}{n}\tilde{\bm{Z}}\bm{Y}=$ $\begin{bmatrix}1\\ (1-s)(2p-1)\\ s\frac{\sigma_{\xi}}{m}\bm{Y}^{\prime}\\ \bm{0}\end{bmatrix}.$ (78)
@@
$\frac{1}{n}(\bm{Z}\bm{Z}^{\prime\prime\top}+\bm{Z}^{\prime\prime}\bm{Z}^{\top})=\begin{bmatrix}\bm{H}&\bm{N}&\bm{0}\\ \bm{N}^{\top}&\bm{0}&\bm{0}\\ \bm{0}&\bm{0}&\bm{0}\\ \end{bmatrix},$ (92)
@@
$=$ $\begin{bmatrix}1&0&\frac{\sigma_{\xi}}{2}&\frac{\sigma_{\xi}}{2}&\dots&-\frac{\sigma_{\xi}}{2}&-\frac{\sigma_{\xi}}{2}&\dots&0&\dots\\ 2p-1&0&(2p-1)\frac{\sigma_{\xi}}{2}&(2p-1)\frac{\sigma_{\xi}}{2}&\dots&-(2p-1)\frac{\sigma_{\xi}}{2}&-(2p-1)\frac{\sigma_{\xi}}{2}&\dots&0&\dots\\ 0&\dots&&&&&&&0&\dots\\ \vdots&&&&&&&&\vdots&\ddots\\ 0&\dots&&&&&&&0&\dots\\ \end{bmatrix}.$ (106)
@@
$\frac{1}{n}\bm{Z}^{\prime\prime}\bm{Z}^{\prime\prime\top}=\begin{bmatrix}\bm{G}_{core}&\frac{\sigma_{\xi}}{m}\bm{F}^{\prime}&\bm{0}\\ \frac{\sigma_{\xi}}{m}\bm{F}^{\prime\top}&\frac{\sigma_{\xi}^{2}}{m}\bm{I}_{m}&\bm{0}\\ \bm{0}&\bm{0}&\bm{0}\end{bmatrix}.$ (110)
@@
$=$ $\frac{1}{m}\begin{bmatrix}\bm{F}^{\prime}\\ \sigma_{\xi}\bm{I}_{m}\\ \bm{0}\end{bmatrix}[\bm{F}^{\prime\top}~{}~{}~{}\sigma_{\xi}\bm{I}_{m}~{}~{}~{}\bm{0}^{\top}]$ (114)
$=$ $\frac{1}{m}\begin{bmatrix}\bm{F}^{\prime}\bm{F}^{\prime\top}&\sigma_{\xi}\bm{F}^{\prime}&\bm{0}\\ \sigma_{\xi}\bm{F}^{\prime\top}&\sigma_{\xi}^{2}\bm{I}_{m}&\bm{0}\\ \bm{0}&\bm{0}&\bm{0}\end{bmatrix}$ (118)
$=$ $\begin{bmatrix}\bm{G}_{core}&\frac{\sigma_{\xi}}{m}\bm{F}^{\prime}&\bm{0}\\ \frac{\sigma_{\xi}}{m}\bm{F}^{\prime\top}&\frac{\sigma_{\xi}^{2}}{m}\bm{I}_{m}&\bm{0}\\ \bm{0}&\bm{0}&\bm{0}\end{bmatrix}$ (122)
@@
$=$ $(1-s)^{2}\begin{bmatrix}\bm{G}_{spu}&\bm{0}\\ \bm{0}&\frac{\sigma_{\xi}^{2}}{d-2}\bm{I}_{d-2}\end{bmatrix}+s(1-s)\begin{bmatrix}\bm{H}&\bm{N}&\bm{0}\\ \bm{N}^{\top}&\bm{0}&\bm{0}\\ \bm{0}&\bm{0}&\bm{0}\\ \end{bmatrix}+s^{2}\begin{bmatrix}\bm{G}_{core}&\frac{\sigma_{\xi}}{m}\bm{F}^{\prime}&\bm{0}\\ \frac{\sigma_{\xi}}{m}\bm{F}^{\prime\top}&\frac{\sigma_{\xi}^{2}}{m}\bm{I}_{m}&\bm{0}\\ \bm{0}&\bm{0}&\bm{0}\end{bmatrix}+\lambda\bm{I}_{d}$ (131)
$=$ $\begin{bmatrix}\bm{A}&\bm{B}&\bm{0}\\ \bm{B}^{\top}&\bm{D}&\bm{0}\\ \bm{0}&\bm{0}&\frac{\sigma_{\xi}^{2}}{d-2}\bm{I}_{d-2-m}+\lambda\bm{I}_{d-2-m}\end{bmatrix},$ (135)
@@
$=$ $\begin{bmatrix}\psi_{1}&\psi_{2}\\ \psi_{2}&\psi_{3}\end{bmatrix},$ (146)
@@
$(\bm{P}/\bm{D})^{-1}=\frac{1}{\psi_{1}\psi_{3}-\psi_{2}^{2}}\begin{bmatrix}\psi_{3}&-\psi_{2}\\ -\psi_{2}&\psi_{1}\end{bmatrix}.$ (149)
Let $\bm{w}$ be the minimizer of the objective in 2, which has the closed form expression $\bm{M}^{-1}\frac{1}{n}\tilde{\bm{Z}}\bm{Y}$. We are ready to derive the elements in $\bm{w}$ using Corollary A.2 and 139. But before that, we first derive the following
@@
$=$ $s(1-s)\sigma_{\xi}\begin{bmatrix}1\\ 2p-1\end{bmatrix}+s^{2}\sigma_{\xi}\begin{bmatrix}1\\ 0\end{bmatrix}$ (154)
$=$ $s\sigma_{\xi}\begin{bmatrix}1\\ (1-s)(2p-1)\end{bmatrix}.$ (157)
Now we can drive the first two elements in $\bm{w}$ with the above and Corollary A.2 and 139
@@
$=$ $q(\bm{P}/\bm{D})^{-1}\begin{bmatrix}1\\ (1-s)(2p-1)\end{bmatrix}$ (163)
$=$ $\frac{q}{\psi_{1}\psi_{3}-\psi_{2}^{2}}\begin{bmatrix}\psi_{3}-\psi_{2}(1-s)(2p-1)\\ -\psi_{2}+\psi_{1}(1-s)(2p-1)\end{bmatrix},$ (166)
@@
$\bm{w}_{3:m+2}=$ $-\bm{D}^{-1}\bm{B}^{\top}(\bm{P}/\bm{D})^{-1}\begin{bmatrix}1\\ (1-s)(2p-1)\end{bmatrix}+\bm{D}^{-1}s\frac{\sigma_{\xi}}{m}\bm{Y}^{\prime}+\bm{D}^{-1}\bm{B}^{\top}(\bm{P}/\bm{D})^{-1}\bm{B}\bm{D}^{-1}s\frac{\sigma_{\xi}}{m}\bm{Y}^{\prime}$ (169)
$=$ $\frac{\sigma_{\xi}}{m}\begin{bmatrix}h_{1}\\ h_{2}\\ \vdots\\ h_{m}\end{bmatrix},$ (174)

### v3

# Few-shot Adaptation to Distribution Shifts By Mixing Source and Target Embeddings

@@ 2 Related Work
Few-shot adaptation and sample efficiency. In scenarios where labeled target data are available but scarce, adapting the last layer to the target distribution in a data-efficient manner presents a significant challenge. The study most relevant to our paper is (Chen et al., 2023), which considers the same setting as ours, where the number of available target data for linear probing is very small (e.g., 2 to 32 per class). They propose Pro2 that first learns a projection that maps the source data embeddings to orthogonal informative directions. Then, it passes the target data embeddings through the learned projection and perform linear probing using these projected embeddings. (Teney et al., 2022)’s method shares a similar intuition but does so with an additional loss term instead of explicitly enforcing orthogonality. However, we find that in both methods, the projection layer learned on the source data may miss directions in the embeddings that are important for the target distribution, as we will theoretically demonstrate. In contrast, our method does not suffer from this issue. We include these two methods as baselines in our experiments. (Zhu et al., 2022) and (Zhang et al., 2022) also utilize both source and target data to achieve few-shot adaptation, albeit in different ways. However, their methods modifies the pretrained encoder, which could lead to overfitting on the limited dataset and degrades representation quality. We provide a comparison and further discussion of these methods in Appendix C.4.
@@ 5.1 A case study of domain generalization: the advantage of MixPro over prior work
(2) MixPro learns $\bm{w}_{MixPro}^{*}=\arg\min_{\bm{w}}\hat{\mathbbm{E}}_{(\bm{z},y)\in\mathcal{E}_{\text{mixed}}}\allowbreak l(\bm{z}^{\top}\bm{w},y)$, as described in Section 4. When $\exists\epsilon=\Theta(1)$ s.t. $\epsilon<s<1-\epsilon$, it achieves a test loss on target data that can be upper bounded by:
@@ 5.2 A case study of subpopulation shift: trade-off between learning target and overfitting noise
Consider the linear model $\bm{w}_{MixPro}^{*}$ learned by MixPro with $l(.,.)$ being MSE loss with $\ell_{2}$ regularization. With assumption 5.2, the test loss on the target distribution achieved by $\bm{w}_{MixPro}^{*}$ can be expressed in closed form in terms of $p_{\text{spu}},\sigma_{1},r$ and $s$, as shown in Appendix A.4.
@@ 6 Experiments
Datasets. We consider 8 datasets with different types of distribution shifts. These include 3 subpopulation shift benchmarks: WaterBirds (Sagawa et al., 2019), bFFHQ (Kim et al., 2021), and UrbanCars (Li et al., 2023); along with 5 domain generalization benchmarks: PACS (Li et al., 2017), VLCS (Fang et al., 2013), Office-Home (Venkateswara et al., 2017), Terra Incognita (Beery et al., 2018), and Camelyon17 (Koh et al., 2021). Due to space limitations, we defer the details to Appendices B.1 and B.2. For the target dataset size, we consider the few-shot scenario where {2, 4, 8, 16} data points per class are sampled from the target distribution (except for Terra Incognita, for which we use only {2, 4, 8} samples, because the smallest class, i.e., class index 9, has only 12 examples). The test set is constructed using the remaining data in the target distribution.
@@ 6.2 Results with cross validation using few-shot data
6.3 Relaxing the requirement of source data availability
In some scenarios, the source and target data may not be available simultaneously. In other words, by the time we encounter the target data and wish to adapt to it, the source data may no longer be accessible. However, MixPro requires both data to be available at the same time to train a linear probe on the mixed source and target embeddings. This implies that one may always need to store the source embeddings to use them when a new target data is encountered. To alleviate this requirement, we propose a simple approach of storing only minimal information about the source data, thereby making the method more flexible.
Specifically, we store only the average embeddings (mean) of each class, resulting in only #classes embeddings. When we create the mixed embeddings, we combine each target data point with the corresponding class mean from the source. We call this modified method MixPro-CM and test it on all the datasets under the same settings as shown in Figure 4. From the figure, we observe that MixPro-CM performs slightly worse than the original MixPro, but still outperforms all other methods. We believe that further exploration into how to improve it, such as retaining more fine-grained information than just class means, would be a valuable direction for future research.
@@ 7 Conclusion
Acknowledgment.
This research was supported by the National Science Foundation CAREER Award 2146492 and Cisco Systems.
@@ A.2.1 Analysis of MixPro
By applying the classical result for the inverse of perturbation (Demmel, 1992) and combining it with 12, we obtain
@@
$\bm{u}_{\text{mixed}}\coloneqq\mathbbm{E}\bm{u}_{\text{mixed}}=\begin{bmatrix}\frac{1}{2}(1-s)\\ \frac{1}{2}\\ \frac{1}{2}s\\ \end{bmatrix}.$
@@
$\bm{\Sigma_{t}}=\begin{bmatrix}\sigma^{2}&0&0\\ 0&\frac{1}{2}&0\\ 0&0&\frac{1}{2}\end{bmatrix}.$
@@ A.2.2 Analysis of Pro2
$=$ $\bm{\Pi}_{i}^{\top}\begin{bmatrix}1/2&0&0\\ 0&1/2&0\\ 0&0&\sigma^{2}\end{bmatrix}\bm{\Pi}_{i}-[1,~{}~{}1,~{}~{}0]~{}~{}\bm{\Pi}_{i}+1.$
@@
$\bm{\Pi}_{i}^{*}=\arg\min_{\bm{\Pi}_{i}}\bigg{(}\bm{\Pi}_{i}^{\top}\begin{bmatrix}1/2&0&0\\ 0&1/2&0\\ 0&0&\sigma^{2}\end{bmatrix}\bm{\Pi}_{i}+1\bigg{)}.$
@@
$\hat{\mathbbm{E}}_{(\bm{z}^{t},y)\in\mathcal{E}^{t}}(\bm{\Pi}^{*\top}\bm{z}^{t}\bm{z}^{t\top}\bm{\Pi}^{*})=\hat{\mathbbm{E}}_{(\bm{z}^{t},y)\in\mathcal{E}^{t}}\begin{bmatrix}(\bm{z}^{t\top}\bm{\Pi}_{1}^{*})^{2}&0&0\\ 0&0&0\\ 0&0&0\end{bmatrix}$
@@ A.3 Effect of $s$ when $\sigma=\omega(1)$
$\mathbbm{E}_{(\bm{z},y)\in\mathcal{E}^{t}}(\bm{w}_{MixPro}^{*\top}\bm{z}-y)^{2}=$ $\bigg{(}\frac{q}{\psi_{1}\psi_{3}-\psi_{2}^{2}}(\psi_{3}-\psi_{2}(1-s)(2p-1))-1\bigg{)}^{2}$
$+$ $\bigg{(}\frac{q}{\psi_{1}\psi_{3}-\psi_{2}^{2}}(-\psi_{2}+\psi_{1}(1-s)(2p-1))\bigg{)}^{2}$
@@
$\frac{1}{n}\bm{Z}\bm{Z}^{\top}=\frac{1}{n}\sum_{i=1}^{n}\bm{z}\bm{z}^{\top}=\begin{bmatrix}\bm{G}_{spu}&\bm{0}\\ \bm{0}&\frac{\sigma_{\xi}^{2}}{d-2}\bm{I}_{d-2}\end{bmatrix},\text{ where }~{}~{}\bm{G}_{spu}\coloneqq\begin{bmatrix}1+\sigma_{1}^{2}&2p-1\\ 2p-1&1\end{bmatrix}$
@@
The following derivations are performed based on the equations in Proposition A.1. We index the examples such that the first $m/2$ examples in $\mathcal{E}^{t}$ have label $1$ and the rest have label $-1$. Since equations 18 and 19 are true, without loss of generality, we assume $\bm{\zeta}^{\prime}_{i}=\sigma_{\xi}\bm{e}_{i}$ where $\bm{e}_{i}$ is the $i$-th standard basis in $\mathbbm{R}^{d-2}$.
@@
$\frac{1}{n}\tilde{\bm{Z}}\bm{Y}=$ $\begin{bmatrix}1\\ (1-s)(2p-1)\\ s\frac{\sigma_{\xi}}{m}\bm{Y}^{\prime}\\ \bm{0}\end{bmatrix}.$
@@
$\frac{1}{n}(\bm{Z}\bm{Z}^{\prime\prime\top}+\bm{Z}^{\prime\prime}\bm{Z}^{\top})=\begin{bmatrix}\bm{H}&\bm{N}&\bm{0}\\ \bm{N}^{\top}&\bm{0}&\bm{0}\\ \bm{0}&\bm{0}&\bm{0}\\ \end{bmatrix},$
@@
$=$ $\begin{bmatrix}1&0&\frac{\sigma_{\xi}}{2}&\frac{\sigma_{\xi}}{2}&\dots&-\frac{\sigma_{\xi}}{2}&-\frac{\sigma_{\xi}}{2}&\dots&0&\dots\\ 2p-1&0&(2p-1)\frac{\sigma_{\xi}}{2}&(2p-1)\frac{\sigma_{\xi}}{2}&\dots&-(2p-1)\frac{\sigma_{\xi}}{2}&-(2p-1)\frac{\sigma_{\xi}}{2}&\dots&0&\dots\\ 0&\dots&&&&&&&0&\dots\\ \vdots&&&&&&&&\vdots&\ddots\\ 0&\dots&&&&&&&0&\dots\\ \end{bmatrix}.$
@@
$\frac{1}{n}\bm{Z}^{\prime\prime}\bm{Z}^{\prime\prime\top}=\begin{bmatrix}\bm{G}_{core}&\frac{\sigma_{\xi}}{m}\bm{F}^{\prime}&\bm{0}\\ \frac{\sigma_{\xi}}{m}\bm{F}^{\prime\top}&\frac{\sigma_{\xi}^{2}}{m}\bm{I}_{m}&\bm{0}\\ \bm{0}&\bm{0}&\bm{0}\end{bmatrix}.$
@@
$=$ $\frac{1}{m}\begin{bmatrix}\bm{F}^{\prime}\\ \sigma_{\xi}\bm{I}_{m}\\ \bm{0}\end{bmatrix}[\bm{F}^{\prime\top}~{}~{}~{}\sigma_{\xi}\bm{I}_{m}~{}~{}~{}\bm{0}^{\top}]$
$=$ $\frac{1}{m}\begin{bmatrix}\bm{F}^{\prime}\bm{F}^{\prime\top}&\sigma_{\xi}\bm{F}^{\prime}&\bm{0}\\ \sigma_{\xi}\bm{F}^{\prime\top}&\sigma_{\xi}^{2}\bm{I}_{m}&\bm{0}\\ \bm{0}&\bm{0}&\bm{0}\end{bmatrix}$
$=$ $\begin{bmatrix}\bm{G}_{core}&\frac{\sigma_{\xi}}{m}\bm{F}^{\prime}&\bm{0}\\ \frac{\sigma_{\xi}}{m}\bm{F}^{\prime\top}&\frac{\sigma_{\xi}^{2}}{m}\bm{I}_{m}&\bm{0}\\ \bm{0}&\bm{0}&\bm{0}\end{bmatrix}$
@@
$=$ $(1-s)^{2}\begin{bmatrix}\bm{G}_{spu}&\bm{0}\\ \bm{0}&\frac{\sigma_{\xi}^{2}}{d-2}\bm{I}_{d-2}\end{bmatrix}+s(1-s)\begin{bmatrix}\bm{H}&\bm{N}&\bm{0}\\ \bm{N}^{\top}&\bm{0}&\bm{0}\\ \bm{0}&\bm{0}&\bm{0}\\ \end{bmatrix}+s^{2}\begin{bmatrix}\bm{G}_{core}&\frac{\sigma_{\xi}}{m}\bm{F}^{\prime}&\bm{0}\\ \frac{\sigma_{\xi}}{m}\bm{F}^{\prime\top}&\frac{\sigma_{\xi}^{2}}{m}\bm{I}_{m}&\bm{0}\\ \bm{0}&\bm{0}&\bm{0}\end{bmatrix}+\lambda\bm{I}_{d}$
$=$ $\begin{bmatrix}\bm{A}&\bm{B}&\bm{0}\\ \bm{B}^{\top}&\bm{D}&\bm{0}\\ \bm{0}&\bm{0}&\frac{\sigma_{\xi}^{2}}{d-2}\bm{I}_{d-2-m}+\lambda\bm{I}_{d-2-m}\end{bmatrix},$
@@
$=$ $\begin{bmatrix}\psi_{1}&\psi_{2}\\ \psi_{2}&\psi_{3}\end{bmatrix},$
@@
$(\bm{P}/\bm{D})^{-1}=\frac{1}{\psi_{1}\psi_{3}-\psi_{2}^{2}}\begin{bmatrix}\psi_{3}&-\psi_{2}\\ -\psi_{2}&\psi_{1}\end{bmatrix}.$
Let $\bm{w}$ be the minimizer of the objective in 2, which has the closed form expression $\bm{M}^{-1}\frac{1}{n}\tilde{\bm{Z}}\bm{Y}$. We are ready to derive the elements in $\bm{w}$ using Corollary A.2 and 26. But before that, we first derive the following
@@
$=$ $s(1-s)\sigma_{\xi}\begin{bmatrix}1\\ 2p-1\end{bmatrix}+s^{2}\sigma_{\xi}\begin{bmatrix}1\\ 0\end{bmatrix}$
$=$ $s\sigma_{\xi}\begin{bmatrix}1\\ (1-s)(2p-1)\end{bmatrix}.$
Now we can drive the first two elements in $\bm{w}$ with the above and Corollary A.2 and 26
@@
$=$ $q(\bm{P}/\bm{D})^{-1}\begin{bmatrix}1\\ (1-s)(2p-1)\end{bmatrix}$
$=$ $\frac{q}{\psi_{1}\psi_{3}-\psi_{2}^{2}}\begin{bmatrix}\psi_{3}-\psi_{2}(1-s)(2p-1)\\ -\psi_{2}+\psi_{1}(1-s)(2p-1)\end{bmatrix},$
@@
$\bm{w}_{3:m+2}=$ $-\bm{D}^{-1}\bm{B}^{\top}(\bm{P}/\bm{D})^{-1}\begin{bmatrix}1\\ (1-s)(2p-1)\end{bmatrix}+\bm{D}^{-1}s\frac{\sigma_{\xi}}{m}\bm{Y}^{\prime}+\bm{D}^{-1}\bm{B}^{\top}(\bm{P}/\bm{D})^{-1}\bm{B}\bm{D}^{-1}s\frac{\sigma_{\xi}}{m}\bm{Y}^{\prime}$
$=$ $\frac{\sigma_{\xi}}{m}\begin{bmatrix}h_{1}\\ h_{2}\\ \vdots\\ h_{m}\end{bmatrix},$